OpenStudy (anonymous):

find exact value of tan 17pi/4

geerky42 (geerky42):

HINT: \(\dfrac{17\pi}{4} = 4\left(\dfrac{4\pi}{4}\right) + \dfrac{\pi}{4}\)

geerky42 (geerky42):

So, \(\tan\dfrac{17\pi}{4} = \tan\dfrac{\pi}{4}\)

geerky42 (geerky42):

Does this helps you?

OpenStudy (anonymous):

yes thank you.